The table below shows the daily production of sweatshirts at the CoolPrintz Screenprinting firm. Diminishing marginal returns to labor set in after the worker is hired. Workers Output 12 30 45 50 54 56 56 54 4 6 7 A. 2nd B. 3rd C. 5th D. 8th
Answer #1
Workers Marginal Return
1 12
2 30-12 = 18
3

45 - 30 =15

4 50-45 = 5
5 54-50 = 4
6 56-54 = 2
7 56-56 = 0
8 54- 56 = -2

Hence, diminishing returns sets in after 8 labors are hired.

Thanks!